OpenStudy (anonymous):

Calculate the force at sea level that a boy of mass 50 kg exerts on a chair in which he is sitting.

OpenStudy (anonymous):

Force = mass x acceleration = 50 kg x 10 m/s^2 = 500 N

OpenStudy (anonymous):

How did you get 10m/s^2 ?

OpenStudy (anonymous):

that is value of g i.e. acceleration due to gravity

OpenStudy (anonymous):

for quick calculation we take g= 10 m/s^2 but for more accurate calculation you can use g=9.8 m/s^2, then force = 50 kg x 9.8 m/s^2 = 490 N

OpenStudy (anonymous):

Ohh. Okay, I see now. Thanks :)

OpenStudy (anonymous):

u r welcome...☺